1) During the months of January, Pearson Corporation sold goods to customers. Assume Pearson uses a...

1)

During the months of January, Pearson Corporation sold goods to customers. Assume Pearson uses a perpetual inventory system. The sequence of events was as follows:

Date

Transaction

Jan.6

Sold goods for $1,290 to Kate Inc. with terms 1/15, n/30. The goods cost Pearson $467.

Jan 10

Sold goods to Randall Corp for $1,560 with terms 2/15, n/30. The goods cost Pearson $746.

Jan 19

Collected cash due from Randall Inc.

What is the gross profit from these transactions?

2) ABC Company sells a product and a 12-month service package for that for a combined price of $980. Separately, the product and the service sell for $300 and $900, respectively. How much of the combined price should be allocated to the product?

Please show all work/calculations, thank you.
Answer #1

1) Calculate gross profit

Sales revenue (1290+1560) 2850
Less: Sales discount (1560*2%) -31.20
Net sales 2818.80
Less: Cost of goods sold (467+746) 1213
Gross profit 1605.80

2) Allocation:

Product allocated cost = 980/1200*300 = 245

Service allocated cost = 980*900/1200 = 735